Klapperich Lab:Notebook/Lab Meeting Notes/2010/02/04

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

4 February 2010 Lab Meeting

Qingqing Prospectus Practice Presentation.

‡ Announcements

  • Oakridge deadline is Confusing. The website says something different from the email. Have one from QQ- Anyone else?
  • Shichu Huang is here - WELCOME!

‡ Flu R01:Integration

  • meetings every other M 1-3pm. 705.

† HDA (Lead: Jaephil, Team:Sonali)

  • Start planning R01 for Submission on 6/5/10. MM, JD, CMK.
  • HDA mix in dried form - Cathie found it
  • LOD tests HDA in a cup - barely amplified at 3pg/mL
  • HDA negative water reactions all have primer dimers/non-specific band at 70bp for c diff HDA (MM)
  • 2nd HDA design - "microSPE" + "Flu primer" + "HDA in a cup" + "Detection kit from Biohelix"

† Sample Concentration (Lead: Jane, Team: Jaephil)

  • New design transfer to COP
  • Air resistance in the air channel is too large, redesign channel to reduce air resistance: now testing with cutter plotter
  • Leakage through the air channel from the fluid channel: now moving air pots farther from fluid channel
  • Set up COMSOL and StarCD, look for governing equations for simulation of evaporation: registered for COMSOL course Feb 25

† SPE Column Optimization for DNA/RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)

  • Sonali put out a new RNA protocol (correct SPE batter, homemade buffers, less ethanol washing, RNA Secure) before break.
  • Protocol tested with 10ng, 1ng, 0.1ng Stratagene purified human RNA (MM) and 100 uL of 1:64 dilution ATCC MDCK sup Flu A/PR/8/34 (JC)
  • Human Total RNA see % yields at 50% and fraction 1 + fraction 2 = 4-5ng RNA out of 10ng with all changes. With high vol *ethanol wash but other protocol improvements, see 40% yield at all other concentrations.
  • For Flu virus see improved Ct value by 7 cycles in fraction 1 from multiple channels
  • Also detecting 1.4 logs more of virus at 1:64 JC virus prep than before.
  • Switching from lambda phage DNA to TaqMan RNase P detection [update]
  • Hussam Runs chip with RNA (confirm procedure is correct) [update]
  • Jessie will run chips with serial virus dilutions from 1:1 to 1:1,000,000 to re-do previous dilution experimnent with WHO SYBR green assay on chip.
  • Jessie will test 2mm wide channels with 4uL SPE with Black Beauty for efficiency.
  • Sonali will assay BIDMC virus sample with newest protocol with 0.7uM Silica and compare with Qiagen RNA extraction with CDC PCR assay

† integrated chip for flu(SPE+RT+PCR) (Lead: Qingqing)

  • integrated chip works with ATCC flu virus
  • working on improving the efficiency
  • testing real patient sample
  • PCR1 Paper back to MCK this week.
  • CMK: PCR 1 draft. Analytical Chem.

‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • See email notes.

† PCR of C.Difficile DNA (Qingqing)

  • ATCC C.dif DNA with Negative patient sample works on chip
  • Another positive patient sample was tested on chip
 -wrong size products(55bp).
 -no designed size product.

‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • New meeting time, T, 9:30 am every other week.
  • Will be done by next Tuesday - PCR with 3 different kits of same serial dilution RNA sample (Invitrogen - SuperscriptIII Platinum, Ambion - AgPath-ID, Qiagen - OneStep RT-PCR Kit)

‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paper on HotDog in progress.

‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • Evap paper - Additional experiment for Fig 3.

‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
* Cathie will submit paper inquiry.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • First samples collected.

‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))

  • Submitted Gantt Document to PATH 11/25.
  • Frank: Repeating 1-week stabilized extractions today (4 Feb).
  • Sean: Started making straws for the new experiments we talked about (DNA weights of 1ng, 10ng, and 100ng through 700nm at SPE 100 monolith at 0-day time point). Have contacted the physics department shop floor to check the progress of the Blackbird pieces I ordered on Monday.
  • Mark: as of 4 Feb. has loaded 6 and 4 month trials with 10ng/straw. Results of the 1ng/straw 0-day trial inconclusive, redo possible. Sticking with 10ng/straw for remaining tests. Loaded 2 month trial on Tuesday. Will load shorter time periods next week (i.e. 3-day, 1 week). Straw fabrication ongoing.

‡ IIH Senior Project.

  • ELISA kits have been delivered
  • Protocol written for using cutter plotter. Will post to share drive
  • Made "test" chips from 100 micron COP but cuts are sloppy
    • Will change settings on cutter to clean up cuts
      • First thought is to slow down speed
  • Started acquiring images of the 100 micron "test" chips